OpenStudy (anonymous):

a 2 kilo gram object falls 3 meters. The potential energy is 58.8, how much work was accomplished by the fall?

OpenStudy (anonymous):

Gravity did 58.8 J work on the object, and the object did not do any work itself, just gained kinetic energy.

OpenStudy (loser66):

\(P = \dfrac{W}{t}\rightarrow W = Pt\) we have P = 58.8 , so just calculate t Since the object is free fall , so, Potential energy + Kinetic energy at initial position = Potential energy + Kinetic energy at the end. \(PE_i +KE_i = PE_f+KE_f\) \( 58.8 + 0 = 0 +\dfrac{1}{2} mv^2\) \(\rightarrow v = \sqrt{\dfrac{2*58.8}{2kg}}=7.67m/s\) from this v, we can calculate t by \(v=v_0 + gt\rightarrow t = \dfrac{v}{g}=\dfrac{7.67}{9.8}=0.78seconds\) now plug back to W = Pt = 58.8 * 0.78 = 46 J

OpenStudy (anonymous):

@Loser66 I think you used the equation for Power, not potential energy!! \[W=-\Delta PE=-mg\Delta h = mg(h_1-h_2)=(2kg)(9.8 \ m/s^2)(3 m)\]=58.8J like @douglaswinslowcooper said ^_^

OpenStudy (loser66):

@AIITehMaffs If you said so, I have nothing to say, but it doesn't make sense to me. XD